OpenStudy (sparklestaraa):

Arnold deposited some money in a bank. The bank paid him a simple interest every year. After t years, the amount of money f(t), in dollars, that the bank paid Arnold was $1,000 more than 20 times the number of years. Which statement best describes the steps to sketch a graph showing the relationship between f(t) and t?

OpenStudy (sparklestaraa):

A. Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 20 and y-intercept = 1000. B. Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 1000 and y-intercept = 20. C. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 20 and y-intercept = 50. D. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 1000 and y-intercept = 50. @undeadknight26

OpenStudy (sparklestaraa):

@CGGURUMANJUNATH

OpenStudy (sparklestaraa):

@satellite73

OpenStudy (sparklestaraa):

@Directrix I think it is B

Directrix (directrix):

Did you get the equation first?

Directrix (directrix):

Wrong: B. Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 1000 and y-intercept = 20. This is why you need to write the equation of the relationship first.

OpenStudy (sparklestaraa):

Then its A? @Directrix

Directrix (directrix):

Until you write the equation based on the given info in the problem, I don't know how to get an answer.

OpenStudy (sparklestaraa):

f(t)= 20*y+1,000 ?

Directrix (directrix):

f(t) = 20*t + 1000 according to the problem. Think of: y = m*x + b where m is the slope and b is the yintercept. Here, the slope is 20 and the yintercept is 1000

Directrix (directrix):

Okay, now select an option.

Directrix (directrix):

>>>Then its A? Yes, that is what I got.

Directrix (directrix):

@Sparklestaraa

OpenStudy (sparklestaraa):

:D thanksss :D

OpenStudy (sparklestaraa):

Omg if u say my medal again xD I'm teasing :P